1. Write


(2−i) (4−3i)2


3+2i


(4)


in the form a+bi, where a,b ∈ R.

1
Expert's answer
2022-06-02T16:38:33-0400
"\dfrac{(2−i) (4−3i)^2}{3+2i}=\dfrac{(2−i) (16-24i-9)(3-2i)}{(3)^2+(2)^2}"

"=\dfrac{(6-4i-3i-2) (7-24i)}{13}"

"=\dfrac{(4-7i) (7-24i)}{13}=\dfrac{28-96i-49i-168}{13}"

"=-\dfrac{140}{13}-\dfrac{145}{13}i"